# Solving Inverse Problems in Protein Space Using Diffusion-Based Priors

### v1

@@ Appendix B Optimization parameters
$\lambda_{m}=0.1,\quad\lambda_{s}=\begin{cases}0\text{ if epoch}<\text{3,000}\\ 1\times 10^{-5}\text{ otherwise}\end{cases},\quad\lambda_{d}=3\times 10^{-5}\cdot(r_{t}/r_{0})^{3},$ (20)
and all the momenta are set to 0.9. $r_{t}$ is set to 15 Å for the first 3,000 epochs and linearly decreases to 5 Å during the last 1,000 epochs. The side chain angles ($\mathbf{\chi}$) are sampled every 100 epochs. The time schedule is
@@
Appendix C Correlated Diffusion
@@
where $a$ is a free parameter, $\nu=\frac{1}{\sqrt{1-b^{2}}}$ and $b$ is chosen such that the radius of gyration $R_{g}$ scales with the number of residues as $R_{g}\sim rN^{\alpha}$ ($r\approx 2.0~{}\text{\AA}$, $\alpha\approx 0.4$ [63, 27]).
Appendix D Target Structures
@@
Table 1: Information on target structures.
PDB Nb. of residues Resolution Release date Imaging method $\in$ CASP15
7r5b 127 1.77 Å 2023-02-08 X-ray diffraction No
7qum 130 1.50 Å 2023-03-01 X-ray diffraction No
1cfd 148 N/A 1995-12-07 NMR No
7pzt 160 1.84 Å 2022-11-02 X-ray diffraction Yes
1a2f 291 2.10 Å 1998-11-25 X-ray diffraction No
Appendix E Additional Results
E.1 Structure Completion
@@
1cfd (148) RMSD ($\downarrow$) Chroma* 0.42 1.17 3.06 5.84 9.44 14.28 15.87
ADP-3D 0.20 0.59 1.83 4.62 12.76 12.51 15.29
ELBO ($\uparrow$) Chroma* 5.65 5.97 6.49 5.46 8.87 9.22 8.94
ADP-3D 5.22 5.34 7.60 8.00 8.91 8.72 7.82
@@
1a2f (291) RMSD ($\downarrow$) Chroma* 0.39 1.21 3.42 7.45 14.77 16.99 17.29
ADP-3D 0.13 0.68 1.91 5.48 12.48 17.78 19.68
ELBO ($\uparrow$) Chroma* 5.39 5.33 6.83 7.09 8.74 9.28 9.48
ADP-3D 5.24 6.54 7.62 7.85 8.20 8.27 9.12
E.2 Gradient Descent for Linear Constraints
@@
Figure 6: Gradient Descent for Linear Constraint. Convergence speed of different techniques on a linear inverse problem (structure completion with a subsampling factor of 2), without the diffusion prior. Using preconditioning with momentum leads to the fastest convergence. The “loss” corresponds to the negative log-likelihood.
E.3 Ablation Study for Model Refinement
@@
E.4 Additional video

### v2

@@ Appendix B Optimization parameters
$\lambda_{m}=0.1,\quad\lambda_{s}=\begin{cases}0\text{ if epoch}<\text{3,000}\\ 1\times 10^{-5}\text{ otherwise}\end{cases},\quad\lambda_{d}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.01},$ (20)
and all the momenta are set to 0.9. $r_{t}$ is set to 5 Å for the first 3,000 epochs and linearly decreases to 1.5 Å during the last 1,000 epochs. The side chain angles ($\mathbf{\chi}$) are sampled every 100 epochs. The time schedule is
@@
Appendix C Compute Time
In each experiment, we run 8 replicas in parallel on a single NVIDIA A100 GPU. The typical compute time depends on the task, the number of residues in the protein, and the diffusion model. With Chroma, for which the compute time scales sub-quadratically with the number of residues [37], the typical compute time for protein with 100 residues is 2 minutes for the structure completion task, 10 minutes for the “distances to structure” task and 30 minutes for the model refinement task.
Appendix D Correlated Diffusion
Figure 5: Condition number of $\mathbf{R}$ vs. number of residues $N$.
@@ Appendix D Correlated Diffusion
where $a$ is a free parameter, $\nu=\frac{1}{\sqrt{1-b^{2}}}$ and $b$ is chosen such that the radius of gyration $R_{g}$ scales with the number of residues as $R_{g}\sim rN^{\alpha}$ ($r\approx 2.0~{}\text{\AA}$, $\alpha\approx 0.4$ [71, 35]). The condition number of $R$ is shown in Figure 5 as a function of the number of residues.
Appendix E Other Diffusion Models
ADP-3D can be extended to leverage diffusion models that do not operate on Cartesian coordinates. If the diffusion model operates in a Riemannian manifold (like $\text{SE}(3)$), the gradient step on the log-likelihood needs to be generalized using the exponential map ($\hat{\mathbf{z}}_{0}=\exp_{\tilde{\mathbf{z}}_{0}}\left(\sum_{i}\mathbf{v}_{i}\right)$).
For diffusion models that do not use a correlated diffusion process, the correlation matrix $\mathbf{R}$ and the preconditioning strategy described in Section 4.2 can be ignored.
Appendix F Target Structures
@@
Table 1: Information on target structures. We indicate the number of modeled residues (residues included in the atomic model provided in the PDB file) and the number of deposited resides (number of residues in the sequence). We selected all single-chain proteins from the CASP15 database. TBP = to be published.
PDB Modeled Deposited Resolution Release date $\in$ CASP15 Reference
7r5b 127 127 1.77 Å 2023-02-08 No [77]
7qum 130 130 1.50 Å 2023-03-01 No [20]
7pzt 160 160 1.84 Å 2022-11-02 Yes TBP
8em5 92 103 1.95 Å 2023-09-27 Yes [17]
8ok3 121 125 1.50 Å 2023-10-25 Yes [7]
7roa 117 136 1.82 Å 2022-10-12 Yes [16]
8ork 460 460 1.64 Å 2023-12-06 Yes [24]
8c6z 573 630 1.85 Å 2023-05-31 Yes [26]
8dys 409 603 1.80 Å 2022-08-17 Yes TBP
7uww 632 635 1.61 Å 2023-06-14 Yes [56]
8h2n 709 821 4.41 Å 2023-10-11 Yes [8]
7zcx 1040 1424 3.10 Å 2023-06-14 Yes [29]
8sxa 1153 1325 3.30 Å 2024-01-17 Yes [58]
7xkw 499 561 3.10 Å 2022-04-20 No [83]
7wsm 464 520 3.25 Å 2022-01-30 No [85]
7ug0 418 418 2.55 Å 2022-03-23 No [36]
Appendix G Additional Results
G.1 Structure Completion
@@
G.2 Experiments with RFdiffusion
We provide additional results using RFdiffusion [79] as a prior for the “structure completion” and “distances to structure” tasks (Table 3). RFdiffusion describes backbone structures as collections of frames in $\mathbb{R}^{3}\times\text{SO}(3)$. We show that using a prior consistently leads to more accurate structures than just performing gradient descent on the neg-log-likelihood.
Table 3: Results for the “structure completion” and “distances to structure” tasks on additional targets with two different priors (Chroma and RFdiffusion). We report the lowest RMSD over 8 runs and the number of outputs with less than 1.5 Å RMSD between parenthesis. We compare our method to gradient descent on the negative log-likelihood (“No prior”).
PDB Prior Fraction of given coordinates Fraction of given distances
00001/20000 00001/80000 00001/320000 00001/80000 00001/160000
8em5 No prior 6.46 (0) 10.23 (0) 15.79 (0) 11.46 (0) 14.60 (0)
RFdiffusion 1.20 (8) 3.50 (0) 16.45 (0) 1.65 (0) 4.75 (0)
Chroma 0.09 (8) 1.57 (0) 14.26 (0) 1.29 (3) 1.86 (0)
8ok3 No prior 6.17 (0) 8.98 (0) 19.67 (0) 13.11 (0) 14.66 (0)
RFdiffusion 1.16 (8) 3.73 (0) 18.48 (0) 3.87 (0) 9.62 (0)
Chroma 0.10 (6) 1.58 (0) 11.43 (0) 1.10 (5) 1.19 (3)
7roa No prior 6.69 (0) 8.59 (0) 15.16 (0) 13.75 (0) 14.59 (0)
RFdiffusion 0.62 (8) 4.01 (0) 15.04 (0) 2.40 (0) 3.16 (0)
Chroma 0.13 (8) 2.89 (0) 11.57 (0) 1.08 (5) 1.44 (4)
8ork No prior 6.39 (0) 9.70 (0) 15.12 (0) 23.68 (0) 22.41 (0)
RFdiffusion 0.98 (8) 3.47 (0) 16.67 (0) 0.87 (2) 2.29 (0)
Chroma 0.17 (6) 1.76 (0) 15.94 (0) 1.30 (4) 1.34 (3)
8c6z No prior 6.32 (0) 10.20 (0) 17.00 (0) 23.11 (0) 24.13 (0)
RFdiffusion 1.16 (8) 4.13 (0) 17.78 (0) 3.69 (0) 4.12 (0)
Chroma 0.16 (8) 2.40 (0) 15.22 (0) 1.42 (4) 1.46 (2)
8dys No prior 6.08 (0) 9.37 (0) 14.74 (0) 18.72 (0) 19.25 (0)
RFdiffusion 1.23 (8) 4.13 (0) 17.78 (0) 1.61 (0) 2.98 (0)
Chroma 0.22 (3) 2.42 (0) 18.74 (0) 1.65 (0) 1.63 (0)
7uww No prior 6.41 (0) 9.55 (0) 18.14 (0) 25.31 (0) 26.78 (0)
RFdiffusion 1.30 (8) 4.55 (0) 21.24 (0) 1.90 (0) 3.13 (0)
Chroma 0.15 (8) 2.67 (0) 16.12 (0) 1.48 (1) 1.47 (1)
8h2n No prior 6.35 (0) 9.37 (0) 16.57 (0) 26.25 (0) 26.24 (0)
RFdiffusion 1.04 (8) 4.44 (0) 20.30 (0) 4.43 (0) 5.10 (0)
Chroma 0.17 (8) 2.22 (0) 15.59 (0) 1.49 (1) 1.52 (0)
7zcx No prior 6.35 (0) 9.43 (0) 18.14 (0) 41.77 (0) 40.22 (0)
RFdiffusion 1.31 (8) 4.36 (0) 22.95 (0) 6.09 (0) 5.81 (0)
Chroma 0.21 (4) 2.36 (0) 17.33 (0) 1.99 (0) 1.74 (0)
8sxa No prior 6.12 (0) 9.19 (0) 14.32 (0) 34.06 (0) 34.43 (0)
RFdiffusion 0.95 (8) 4.06 (0) 15.70 (0) 1.87 (0) 2.64 (0)
Chroma 0.19 (8) 2.02 (0) 10.78 (0) 1.65 (0) 1.86 (0)
G.3 Gradient Descent for Linear Constraints
@@
Figure 6: Gradient Descent for Linear Constraint. Convergence speed of different optimization techniques on a linear inverse problem (structure completion with a subsampling factor of 2), without the diffusion prior. Using preconditioning with momentum leads to the fastest convergence. The “loss” corresponds to the sum of squared distances between unmasked atom coordinates.
G.4 Comparison to DPS
We adapted the Diffusion Posterior Sampling (DPS) algorithm [12] to tackle the structure completion problem. The central idea of DPS is to combine a reverse diffusion step with a gradient-ascent step on the log-likelihood. Instead of computing the gradient from the iterate $\mathbf{x}_{t}$ (w.r.t. $\mathbf{x}_{t}$), DPS suggests to denoise the current estimate ($\hat{\mathbf{x}}_{0}=\hat{\mathbf{x}}_{\theta}(\mathbf{x}_{t})$) and compute the gradient from the iterate $\hat{\mathbf{x}}_{0}$ (w.r.t. $\mathbf{x}_{t}$). The magnitude of the gradient step is controlled by a parameter named $\zeta^{\prime}$.
We compare ADP-3D to the adaptation of DPS for the structure completion task on PDB:8ok3 with a subsampling factor of 4. We compare the distribution of final RMSDs over 64 replicas and the total runtime. We perform a sweep over $\zeta^{\prime}$ for DPS. Both methods are ran over 1,000 steps. As shown in Figure 7, ADP-3D leads to more accurate reconstructions. Moreover, since ADP-3D does not need to compute gradients through the denoiser, it is approximately six times faster than DPS, per iteration.
Figure 7: Comparison to DPS. We compare ADP-3D to DPS [12] for the structure completion task on PDB:8ok3, with a subsampling factor of 4. (a) Distribution of final RMSDs over 64 replicas. For DPS, we perform a sweep over the parameter $\zeta^{\prime}$, controlling the magnitude of the gradient step. (b) Comparison of runtime for the same experiment.
G.5 Ablation Study for Model Refinement
@@
G.6 Influence of Resolution on Model Refinement
We use ChimeraX [52, 70] to simulate the 3D density maps of the TecA bacterial toxin (PDB:7pzt), at different resolutions. We run ModelAngelo [39] on the simulated maps, using the known sequence and the default parameters. We show our results in Figure 9. For both input resolutions (2.0 Å and 4.0 Å), our method improves on ModelAngelo’s accuracy for a fixed completeness level (and reaches a higher completeness for the same accuracy). As the resolution of the map improves (gets smaller), the completeness of the output of ModelAngelo increases and the output of ADP-3D improves.
Figure 9: Atomic Model Refinement. Results on the TecA bacterial toxin (PDB:7pzt, 160 residues). (a) Qualitative results. From left to right: the input density map at 2.0 Å resolution, the incomplete model given by ModelAngelo and our refined models (1 output and 5 outputs), overlaid on the target structure in transparency. (b) RMSD of alpha carbons vs. completeness (number of predicted residues / total number of residues) with ModelAngelo (MA) and our method. We run 5 experiments and report the mean of the lowest RMSD on $\alpha$-carbons over 8 replicas (±$\pm 1$ std). The spread of RMSD is further described in the supplements. The experimental (deposited) resolution is indicated with a dashed line.
G.7 Variability of the Output on the Model Refinement Task
[…]
In Figure 10, we show the spread of final RMSDs obtained over 40 replicas of the same experiment (refinement of 7pzt with 2 Å density map). Some of the runs fail to reach sub-2Å RMSD, but we show that these outliers can be removed by measuring how well the structures fit into the input density map (using the log-likelihood function). This outlier removal process can therefore operate without access to ground truth information, making it straightforward to apply on new molecules.
Appendix H Additional video